# BiMediX2 : Bio-Medical EXpert LMM for Diverse Medical Modalities

### v1

@@ 4 Results
UPHILL OpenQA Evaluation
Fig 4 illustrates the performance comparison on the UPHILL OpenQA benchmark. BiMediX2 70B achieved the highest overall factual accuracy of 60.6%, and the second highest being BiMediX2 8B (56.1%), surpassing other models such as GPT-4 (51.5%), Meditron 70B (49.6%), and Med42 (53.5%). This highlights BiMediX2’s effectiveness in discerning and correcting misinformation in medical contexts.
@@
Figure 4: Performance comparison on UPHILL OpenQA (Kaur et al. (2023)), assessing the model’s ability to address false medical claims at different presupposition levels.
BiMed-MBench Evaluation
Tables 3 and 4 present the evaluation results of BiMediX2 and other medical LMMs on the English and Arabic BiMed-MBench benchmark, respectively. BiMediX2 8B demonstrated superior performance in both English (overall score of 62.2%) and Arabic (overall score of 50.5%) evaluations, outperforming other models. Fig 1 compares the performance of state-of-the-art medical LMMs on our BiMed-MBench evaluation in a bilingual context. This indicates BiMediX2’s strong bilingual capabilities and its effectiveness in handling medical conversations and descriptions across different imaging modalities.
Medical VQA Benchmark
Tab 5 shows the performance of BiMediX2 and other models on the Medical VQA benchmark using the MultiMedEval toolkit. Our BiMediX2 8B achieves the highest average score of 0.611, outperforming other models across datasets such as Rad-VQA, Slake-VQA, and Path-VQA. This demonstrates BiMediX2’s proficiency in visual question answering, a critical task in medical diagnostics.
@@
Average 0.461 0.239 0.177 0.427 0.319 0.509 0.611
Table 5: Medical VQA Benchmark (MultiMedEval Royer et al. (2024))
Report Summarization
Tab 6 presents the report summarization performance on the MIMIC-III dataset. BiMediX2 8B achieved the highest average score of 0.416, surpassing other models like LLaVA-Med (0.398) and Dragonfly-Med (0.110). This average score is derived as a unified metric by re-scaling BLUE-4* and RadCliQ* metrics. This indicates BiMediX2’s effectiveness in generating concise and accurate summaries of medical reports, a vital task for efficient healthcare communication.
Report Generation
Tab 7 shows the report generation performance on the MIMIC-CXR dataset. BiMediX2 8B achieved the highest average score of 0.235, outperforming other models like LLaVA-Med (0.192) and BioMedGPT (0.145). The average score is derived as a unified metric by re-scaling BLUE-4* and RadCliQ* metrics. This highlights BiMediX2’s capability in generating detailed and accurate medical reports from radiology images, a crucial task for diagnostic purposes.
Dataset Metric LLaVA Med Dragonfly-Med BiMediX2 4B BiMediX2 8B
MIMIC-III ROUGE-L↑ 0.185 0.072 0.209 0.205
BLEU-1↑ 0.192 0.062 0.153 0.178
BLEU-4↑* 0.520 0.000 0.410 0.449
F1-RadGraph↑ 0.232 0.000 0.222 0.230
RadCliQ↑* 0.753 0.247 0.923 0.918
CheXbert vector↑ 0.600 0.326 0.633 0.593
METEOR↑ 0.303 0.060 0.264 0.339
Average 0.398 0.110 0.402 0.416
Table 6: Report Summarization (MultiMedEval Royer et al. (2024))
Dataset Metric RadFM LLaVA Med BioMedGPT BiMediX2 4B BiMediX2 8B
MIMIC-CXR Report Generation F1-RadGraph↑ 0.042 0.048 0.000 0.083 0.098
BLEU-1↑ 0.006 0.163 0.003 0.046 0.155
BLEU-4↑* 0.000 0.060 0.000 0.042 0.016
ROUGE-L↑ 0.065 0.125 0.012 0.131 0.153
RadCliQ↑* 0.655 0.660 0.827 0.865 0.860
CheXbert vector↑ 0.197 0.150 0.153 0.205 0.189
METEOR↑ 0.053 0.137 0.016 0.107 0.174
Average 0.145 0.192 0.145 0.211 0.235
Table 7: Report Generation (MultiMedEval Royer et al. (2024))
Qualitative Examples
Medical Image Understanding in a Conversational Context: Fig 5 illustrates the capabilities of the BiMediX2 framework in analyzing medical images and providing detailed, context-aware responses. The top section showcases a conversation involving a sagittal CT scan of the lumbar spine. The model identifies the scan type and explains that it is a vertical slice of the body focusing on the lumbar spine. When asked about abnormalities, the model accurately identifies a fracture in the L4 vertebra, providing an explanation of the condition and its potential causes, such as trauma or stress. The bottom section features a color Doppler ultrasound scan of the left ovary, identifying a potential abnormality a left ovarian cyst with a solid component. Our model explains the imaging technique, names the organ, and discusses the detected abnormality, emphasizing the need for further evaluation. These examples demonstrate BiMediX2’s ability to interpret complex medical images and provide informative responses to assist in clinical decision-making.
Figure 5: Qualitative Examples of our BiMediX2 for Medical Image Understanding in a Conversational Context.
@@ 6 Limitations
Currently, our models lack explicit mechanisms to curb undesirable behaviors. Future research will focus on enhancing security and alignment strategies. On a brighter note, we believe that releasing our model weights could contribute to investigating and mitigating risks tied to application uses.
@@ 7 Safety and Ethical Implications
We recognize the significant societal impact of BiMediX2 and emphasize the importance of ethical considerations and transparency. This release is intended for research purposes only and is not ready for clinical or commercial use. Ensuring the model’s accuracy and reliability is crucial, as incorrect medical advice could have serious health consequences. Robust validation and quality control measures are essential to minimize errors.
@@ 8 Acknowledgement
We would like to thank Dr. Omair Mohammed, Dr. Mohammed Zidan and Dr. Vishal Thomas Oommen for their contribution in verification of medical responses.
@@
We are grateful to the Meta Llama Impact Innovation Awards for recognizing BiMediX2, as one of the winners awarded in October 2024. This recognition highlights our commitment to advancing AI-driven healthcare solutions.

### v2

@@ 2.1 BiMed-V: Multimodal, Bilingual Dataset
2.2 Medical Instruction Tuning
To enable BiMediX2’s robust capabilities in both bilingual and multimodal medical tasks, we employ a two-stage training process that ensures precise alignment of visual and textual representations while adapting the language model for complex medical instruction tasks (see Fig 2).
Stage 1: Medical Concept Alignment: In the first stage, we finetune the Projector alone to align visual embeddings to the language embedding space. The training utilizes a comprehensive dataset of 467k image-caption pairs sourced from the LLaVA-Med Li et al. (2023) pretraining dataset. These pairs span diverse medical imaging modalities and captions that describe clinically relevant features.
Stage 2: Multimodal Medical Instruction Alignment: The second stage finetunes the LoRAHu et al. (2021) adapters within the language model, enhancing its ability to process and generate multimodal medical instructions. For this, we utilize our BiMed-V 1.6M bilingual multimodal instruction set, which comprises carefully crafted English and Arabic prompts paired with corresponding visual and textual responses. This dataset enables the model to learn nuanced instructions across a wide array of medical domains, from radiology to pathology, in a bilingual context.
Through these two stages, BiMediX2 achieves seamless integration of bilingual and multimodal capabilities, enabling it to deliver accurate and context-aware medical insights in both English and Arabic, tailored to a variety of clinical scenarios. We have used LLaMA 3.1 (8B, 70B) and Phi-3.5 V Abdin et al. (2024) as base models to obtain BiMediX2 8B, BiMediX2 70B, and BiMediX2 4B variants, respectively. For further details on model and training configurations, see Section A.6.
3 Experiments
In the literature, evaluating medical language models predominantly involves multiple-choice question-answering tasks, with accuracy as the performance metric. We employed the EleutherAI Gao et al. (2024) evaluation framework for evaluating text-only medical tasks following Pieri et al. (2024). For our bilingual and multimodal evaluations, we utilized several benchmarks and tools. We leveraged our BiMed-MBench for assessing the correctness, helpfulness, relevance and accuracy of the model in a bilingual medical multimodal context. Additionally, we evaluated the model on VQA using datasets such as Path-VQA, SLAKE, and Rad-VQA. For report generation, we used the MIMIC-CXR Johnson et al. (2019) dataset, and for report summarization, we used the MIMIC-III Johnson et al. (2016) dataset. These evaluations were conducted using the MultiMedEval Royer et al. (2024) toolkit, ensuring a comprehensive assessment of the model’s multimodal capabilities.
3.1 Medical LLM Benchmarks
We assess our models on text-based clinical benchmarks using PubMedQA, MedMCQA, MedQA, USMLE, Medical MMLU, and UPHILL OpenQA. For further details, see Section A.3.
Figure 3: Qualitative Examples of BiMediX2 for Medical Image Understanding in a Conversational Context.
@@ 3.1 Medical LLM Benchmarks
Figure 4: Performance comparison on UPHILL OpenQA Kaur et al. (2023), assessing the model’s ability to address false medical claims at different presupposition levels.
3.2 Medical VLM Benchmarks
BiMed-MBench: We introduced the first bilingual GPT-4o-based medical LMM benchmark, consisting of 386 medical queries spanning various medical imaging modalities. The test dataset includes conversational interactions and detailed descriptions for modalities such as chest X-rays (CXR), MRI, histology, gross and CT scans derived from LLaVA-Med Li et al. (2023). GPT-4o evaluates the correctness of model responses based on the provided image context and caption. The reference prediction, serving as the upper-bound answer for the teacher model, is generated using GPT-4 Achiam et al. (2023). For Arabic evaluations, the ground truth is derived from the upper-bound, translated by GPT-4o, and all test samples are comprehensively verified by medical experts for quality assurance. GPT-4o evaluates the responses by comparing the candidate LMM against the upper-bound for English queries and against the Arabic ground truth for Arabic queries. The evaluation assesses the helpfulness, relevance, accuracy, and level of detail of the responses while maintaining the user language. The scoring is conducted on a scale of 1 to 10, with higher scores indicating better performance. Additionally, GPT-4o provides detailed explanations of the evaluation to offer deeper insights into the models’ performance. Finally, we compute and compare the relative scores of the two candidates.
Visual Question Answering (VQA): To evaluate a model’s performance in Visual Question Answering, three datasets are used: Path-VQA, SLAKE, and Rad-VQA. These datasets include a mix of open-ended and close-ended questions, designed to test the model’s ability to interpret and respond to queries based on medical images. The evaluation employs accuracy, precision, recall, and F1 score to measure the correctness and completeness of the model’s answers, which are critical for clinical reliability. Additionally, BLEU ensures the generated answers are naturally phrased and easy to understand for clinical communication.
@@ 3.2 Medical VLM Benchmarks
[…]
Table 5: Medical VQA Benchmark MultiMedEval Royer et al. (2024)
Report Generation: The MIMIC-CXR Johnson et al. (2019) dataset, which includes de-identified radiology reports associated with chest X-rays, is utilized to assess the model’s performance in generating the findings section of medical reports. For this task, the input consists of one or more radiology images related to a case, followed by a prompt such as, “<<image>> <<image>> Please caption this scan with findings and impressions." This setup evaluates the model’s ability to generate accurate and coherent medical reports based on visual inputs.
Report Summarization: We conduct an evaluation of report summarization using the MIMIC-III Johnson et al. (2016) dataset. The task involves generating the impressions section of a radiology report based on the findings section. To achieve this, the model is provided with the free-text findings along with a task prompt, “Summarize the findings". This assessment focuses on the model’s ability to distill comprehensive medical information into a concise and precise summary, which is essential for clear and effective medical communication.
In both report generation and summarization tasks, relying solely on lexical metrics such as BLEU, ROUGE-L, and METEOR is insufficient, as these do not guarantee clinical accuracy. Therefore, we complement them with clinically-informed metrics such as F1-RadGraph, CheXbert similarity, and RadCliQ. These metrics assess the preservation of key medical entities, relations, and overall clinical correctness, ensuring that the generated or summarized content aligns not just in phrasing but also in diagnostic intent and clinical relevance.
Dataset Metric LLaVA Med Dragonfly-Med BiMediX2 4B BiMediX2 8B Med-PaLM M
MIMIC-III ROUGE-L↑ 0.185 0.072 0.209 0.205 0.320
BLEU-1↑ 0.192 0.062 0.153 0.178 0.154
BLEU-4↑* 0.520 0.000 0.410 0.449 -
F1-RadGraph↑ 0.232 0.000 0.222 0.230 0.347
RadCliQ↑* 0.753 0.247 0.923 0.918 -
CheXbert vector↑ 0.600 0.326 0.633 0.593 -
METEOR↑ 0.303 0.060 0.264 0.339 -
[…]
Table 6: Report Summarization MultiMedEval Royer et al. (2024)
(i) Baseline: VLM trained on non-medical Data. (ii) Stage-1: Aligning Medical images to the LLM input token space by learning an alignment layer and freezing the Language model and the Vision encoder using our pre-training dataset of 467k instructions. (iii) Stage-2: LoRA finetuning with our BiMed-V1.6M dataset comprising a text and image+text dataset across medical image modalities.
Here the Baseline model is obtained by pretraining the projector on LCS-558K dataset following the LLaVA-pp Rasheed et al. (2024) repository. The baseline model lacks understanding of medical images which is introduced in our Stage-1 alignment training. Finally we LoRA finetune model to obtain our medically instruction tuned model. As observed in Tab 9, the text based medical evaluation scores do not change with Stage 1 training as we are only training the projector here and the language model performance is therefore consistent.
Evaluation of English Model with Arabic Translation Pipeline: To further validate the necessity of bilingual instruction tuning, we employed a cascaded translation pipeline for evaluating our English instruction-tuned model (BiMediX2-8B ENG) on the BiMed-MBench Arabic benchmark using popular translation services such as from Google, Alibaba, and Bing.
While this method yielded slight improvements over the base English model as shown in Tab 8, it consistently fell short compared to our bilingual model. These results demonstrate that translation pipelines alone are inadequate for generating medically accurate Arabic responses, highlighting the importance of dedicated bilingual training and instruction tuning. Further experiments are detailed in Section A.5.
Model BiMed-MBench (Ara)
BiMediX2-8B ENG 44.61
BiMediX2-8B ENG
+ Translation (Google)
44.92
BiMediX2-8B ENG
+ Translation (Bing)
45.05
BiMediX2-8B ENG
+ Translation (Alibaba)
45.33
BiMediX2-8B BI 50.47
Table 8: Comparison of Translation frameworks
6 Qualitative Examples
Medical Image Understanding in a Conversational Context: Fig 3 illustrates the capabilities of the BiMediX2 framework in analyzing medical images and providing detailed, context-aware responses in both English and Arabic. The top section highlights BiMediX2 analyzing a sagittal CT scan of the lumbar spine, correctly identifying the scan type and diagnosing an L4 vertebral fracture. It explains potential causes like trauma or stress and discusses clinical implications. In the bottom section, BiMediX2 accurately identifies a female reproductive organ, explains the imaging technique, and detects abnormalities, demonstrating bilingual capabilities in English and Arabic. Additional qualitative examples are provided in Section A.1, while Section A.2 presents its limitations.
@@ 7 Conclusion
Evaluation Baseline Stage-1 Stage-2
BiMed-MBench [3] 34.9 47.6 62.2
Clinical LLM Eval [2] 67.6 67.6 70.4
Table 9: Multi-stage Training Ablation
@@ 8 Limitations
Currently, our models lack explicit mechanisms to curb undesirable behaviors. Future work will focus on enhancing alignment and safety strategies to reduce risks associated with clinical deployment. On a brighter note, we believe that releasing our model weights could contribute to investigating and mitigating these risks through broader community engagement. In addition, our current model does not explicitly mitigate biases related to gender, ethnicity, or socioeconomic status in medical contexts. We recognize the critical importance of fairness in AI-driven healthcare solutions, and acknowledge that such biases could perpetuate or exacerbate disparities in medical decision-making. As part of future work, we plan to conduct targeted bias assessments and explore debiasing strategies. We also believe that open-sourcing our model will facilitate further research into bias detection and mitigation.
Another current limitation is that our model is trained exclusively on Modern Standard Arabic (MSA). This restricts its applicability in real-world clinical settings where dialectal Arabic, such as Emirati, Saudi, Egyptian, or Moroccan is more commonly used. In future iterations of this work, we aim to expand our training corpus to include a diverse range of Arabic dialects, thereby enhancing the model’s utility and inclusivity across Arabic-speaking populations.
@@ 9 Safety and Ethical Implications
We recognize the significant societal impact of BiMediX2 and emphasize the importance of ethical considerations and transparency. This work is intended for research purposes only and is not ready for clinical or commercial use. Ensuring the model’s accuracy and reliability is crucial, as incorrect medical advice could have serious health consequences. Robust validation and quality control measures are essential to minimize errors.
@@ 10 Acknowledgement
We would like to thank Dr. Omair Mohammed, Dr. Mohammed Zidan, Dr. Vishal Thomas Oommen and Dr. Mehrubin Kurpat for their contribution in verification of medical responses.
@@
This work is partially supported by the Meta Llama Impact Innovation Award, the Meta Regional Research Grant, the Google Research Award, the NVIDIA Academic Grant, and the MBZUAI-WIS Research Grant (P008). These recognitions highlight our commitment to advancing AI-driven healthcare solutions.
Appendix A Appendix
A.1 Additional Qualitative Examples
Medical Image Understanding in a Conversational Context: Fig 8 illustrates the capabilities of the BiMediX2 framework in analyzing medical images and providing detailed, context-aware responses in both English and Arabic. The top section showcases a conversation involving a 3D CT scan of the chest. The model identifies the scan type and explains that it uses X-ray technology to create detailed cross-sectional images, which are then reconstructed into 3D images. When asked about abnormalities, the model accurately identifies multiple rib fractures. It further clarifies that these fractures are present on both the left and right sides of the chest. And in the bottom section, BiMediX2 accurately identifies the organ, explains the imaging technique, and detects abnormalities, providing valuable insights that can aid in the diagnosis and treatment of conditions related to the female reproductive system. This particular example showcases BiMediX2’s capability to converse in both English and Arabic simultaneously, depending on the input query.
Figure 5: Data Translation Framework
A.2 Model Failure Cases
We present a common failure case observed in both English and Arabic settings. The example is shown in Fig 7.
In the English-language evaluation, a common source of error was the model’s difficulty in distinguishing between visually similar but clinically distinct fracture types. As shown in Fig 7, while the model response is fluent and medically coherent, it contains a key factual error: the misclassification of an avulsion fracture as a burst fracture. Notably, this response was initially accepted by several medical professionals due to the similarity in language and presentation. However, a trained radiologist later identified the subtle radiographic distinctions, confirming the ground truth diagnosis as an avulsion fracture. This highlights the model’s challenge in handling nuanced diagnostic distinctions that require expert-level domain knowledge.
In the Arabic benchmark, one of the most frequent error modes involved incorrect localization of the affected area as illustrated in Fig 7. The response demonstrates a localization mismatch where the model incorrectly identifies C2–C3 as the affected vertebrae instead of the correct C6–C7 levels. Such errors, especially in high-stakes medical applications, can have serious clinical implications, reinforcing the importance of precise anatomical grounding in medical AI systems.
A.3 Clinical LLM Benchmarks
PubMedQA Jin et al. (2019) is a question-answering dataset derived from biomedical research papers on PubMed. The task involves answering ’yes’, ’no’, or ’maybe’ based on question sourced from the title of a research paper and a context from the abstract. Our analysis focuses on the PQA-L subset, which includes 500 manually annotated QA pairs requiring in-depth biomedical reasoning.
MedMCQA Pal et al. (2022) consists of 4,183 multiple-choice questions from Indian AIIMS and NEET PG medical entrance exams, assessing professional medical knowledge and language comprehension.
MedQA Jin et al. (2021) features multiple-choice questions from medical board exams in the US, Mainland China, and Taiwan. Our study focuses on the USMLE portion (1,273 test samples), requiring multi-step reasoning and evidence retrieval.
USMLE Han et al. (2023) is a self-assessment test from the United States Medical Licensing Examination (Step 1, Step 2, and Step 3). We use the MedAlpaca version, which excludes image-based questions and contains 325 test samples.
Medical MMLU Hendrycks et al. (2020) is a collection of six datasets covering 1,089 test questions across Clinical Knowledge, College Biology, College Medicine, Medical Genetics, Professional Medicine, and Anatomy.
UPHILL OpenQA Kaur et al. (2023) evaluates LLMs’ accuracy in handling health-related queries with varying presuppositions. Our analysis focuses on zero-shot models’ ability to refute false health claims, a critical factor in combating misinformation. In this context, the accuracy refers to the model’s effectiveness in accurately refuting false health-related claims at different presupposition levels.
Fig  6 compares state-of-the-art medical LLMs and LMMs on clinical benchmarks. BiMediX2 70B achieved the highest average score (84.6%), surpassing GPT-4 (82.9%) and Llama-3-Med42-70B (83.0%), demonstrating its superior performance across diverse medical tasks. The stacked bars illustrate performance across individual datasets, highlighting BiMediX2’s strong understanding of medical contexts.
A.4 Data Translation Framework
To construct a high-quality bilingual dataset, we develop a robust data translation framework that leverages GPT-4o for translating English medical content into Arabic, followed by expert human verification to ensure contextual and terminological accuracy. This approach supports effective training and benchmarking in a bilingual context.
We first translate our English multimodal instruction set to Arabic using GPT-4o. A random subset of this translated data is passed to Bilingual medical doctors to verify the quality of the translation. Doctors are asked to report the number of samples requiring minor (e.g rewording, formatting, or stylistic edits) / major (e.g incorrect clinical terms or misinterpretations) revision in the Arabic translation. They are required to look for the overall consistency of the translated text and the accuracy of medical terminologies.
For our BiMed-MBench bilingual benchmark doctors were asked to manually correct these inaccuracies in the translated text. On average 22% of the samples required minor corrections and re-formatting, while only 5% of the samples required major corrections in medical terms.
The verification process involved 10 medical experts from three different countries, ensuring representation from both native Arabic and English speakers. These experts specialized in various medical domains, including radiology (MD Radiology), endocrinology, neurosurgery, general practice, histopathology (MS Pathology), and general medicine. To ensure rigorous validation, a multi-reviewer setup was adopted for a randomly selected subset of the benchmark. Each sample was independently reviewed by multiple doctors, which surfaced occasional inconsistencies, particularly in ambiguous or borderline cases.
To resolve such discrepancies, we introduced a structured adjudication protocol. Conflicting evaluations were circulated among the experts, followed by focused discussion sessions in which reviewers examined each other’s rationale. This collaborative refinement process significantly reduced inconsistencies and improved the reliability of the benchmark.
A.5 Additional Experiments
Model BiMed-MBench (Ara)
BiMediX2-8B ENG 44.6
BiMediX2-8B ARA 46.3
BiMediX2-8B BI 50.5
Table 10: Bilingual vs Monolingual model training
Hyperparameter Stage 1 Stage 2
Number of Epochs 1 1
Train Batch Size (per device) 1 4
Gradient Accumulation Steps 1 4
Learning Rate 1.0×10-3 2.0×10-4
Optimizer Adam Adam
Weight Decay 0 0
Warmup Ratio 0.03 0.03
Learning Rate Scheduler Cosine Cosine
Precision bfloat16 bfloat16
PEFT None LoRA
LoRA Rank - 8
LoRA Alpha - 16
Multi-Modal Projector LR - 2.0×10-5
Table 11: Model training Hyperparameters
A.5.1 Human Expert Evaluation
To assess the clinical quality of model outputs, we conducted a human expert evaluation using a blind review setup. For a randomly selected subset of questions from the BiMed-MBench benchmark, responses were generated by three models: BiMediX2, Dragonfly-Med, and LLaVA-Med. The outputs were anonymized and labeled as Model A, Model B, and Model C, with no identifiers provided to the reviewers.
Medical experts were asked to evaluate the responses against the provided ground truth descriptions for each question. The evaluation focused on determining which model produced the most accurate, clinically relevant, and clear explanation of the medical image.
The results demonstrate a strong preference for BiMediX2, which was selected as the best response in 76.9% of the cases. In comparison, Dragonfly-Med was preferred in 15.4% of the cases, and LLaVA-Med in 7.7%.
Evaluation Protocol for Medical Experts: Your task is to evaluate the responses provided by three AI models based on a given medical image description (Ground Truth). Follow these steps to make your selection:
1) Read the Ground Truth: Carefully review the provided description of the medical image. This serves as the reference for an accurate and detailed response.
2) Assess the Model Responses: Examine the three model-generated responses (Model A, Model B, and Model C). Compare their content with the Ground Truth, focusing on the accuracy, completeness, and relevance of the description.
3) Select the Best Response: Choose the model response that best aligns with the Ground Truth in terms of:
- Clinical Accuracy: Does the response correctly describe the key findings in the image?
- Relevance: Does the response stay focused on the specific details highlighted in the Ground Truth?
- Clarity: Is the explanation well-structured and easy to understand
4) Submit Your Choice: After evaluating the responses, select the one that provides the most accurate and comprehensive explanation.
Dataset No. of samples
PubMedQA 210169
MedMCQA 182712
MedQA 20691
Single Turn QA 119879
Multi-Turn Conversation 133134
PMC-MCQ 80000
LLaVA-Med-Subset-to-Conv 11616
Rad-VQA 1796
Slake-VQA 9835
PMC-VQA 80000
Path-VQA 19654
LLaVA-Med-to-QA 163463
MedQA (Ara) 11210
PubMedQA (Ara) 115773
MedMCQA (Ara) 97523
Single Turn QA (Ara) 156254
Multi-Turn Conversation (Ara) 64235
LLaVA-Med-to-QA (Ara) 163463
PMC-MCQ (Ara) 50000
[…]
Table 12: BiMed-V-1.6M Dataset Composition
A.5.2 Comparison of Bilingual vs Monolingual Training
The results in Tab 10 highlight the impact of language setting on model performance for Arabic medical evaluation tasks. The bilingual model (BiMediX2-8B BI), trained on both English and Arabic instruction data, achieves the highest score of 50.5, outperforming both the Arabic-only model (46.3) and the English-only model (44.6) on the Arabic subset of BiMed-MBench.
This demonstrates that bilingual instruction tuning provides complementary knowledge transfer, enabling the model to better generalize in Arabic. The performance gain over the Arabic-only model further emphasizes the value of incorporating English medical knowledge during training, even when the target evaluation is in Arabic.
A.6 Model Configurations
All traning experiments were conducted with 8× AMD Instinct MI200 GPUs (each with 64 GB of VRAM). The training process is structured in two stages: Stage 1 training requires approximately 20 hours, and Stage 2 training requires approximately 32 hours on this setup. The complete set of training hyperparameters are provided in Tab 11.
We also compare the trade-offs between our model sizes along with their corresponding performance and medical accuracy in Tab 13. We compare the Token Throughput measured in (Tokens/second), Mean Time to First Token (TTFT) (ms), GPU VRAM consumed and Medical Accuracy based on our evaluation benchmarks. These benchmarks are run using the vLLM benchmarking suite Kwon et al. (2023) on 4xNVIDIA RTX A6000 (48GB) GPUs.
A.7 BiMedV-1.6M Dataset
The BiMedV-1.6M dataset comprises 1,691,407 samples across English and Arabic, covering text and image-text QA tasks, supporting multilingual and multimodal medical research. The dataset composition is presented in Tab 12, Fig 9.
Model Token Throughput (tok/s) Mean TTFT (ms) GPU VRAM (GB) Medical Accuracy
BiMediX2-4B 263.05 180.5 42 50.5
BiMediX2-8B 187.14 241.9 44 70.4
BiMediX2-70B 67.61 863.6 170 84.6
Table 13: Performance comparison of BiMediX2 variants
Figure 7: Bilingual failure case: Model labels avulsion fracture as burst (English), and misidentifies C2–C3 instead of C6–C7 (Arabic).
Figure 8: Qualitative Examples of BiMediX2 for Medical Image Understanding in a Conversational Context.
Figure 9: BiMed-V-1.6M Dataset Distribution. The sunburst chart illustrates the hierarchical composition of the dataset, categorizing samples by language (English and Arabic), modality (text and image-text), and task type (QA and VQA). The inner rings represent language and modality divisions, while the outer rings detail specific datasets included in BiMedV-1.6M.